OpenStudy (anonymous):

If cot^2x=5/7, then csc^2 is equivalent to a)12/7 b)7/5 c)5sqrt(74)/74 d)sqrt(74)/7

hartnn (hartnn):

use the relation, \(\csc^2x = 1+ \cot^2 x\)

hartnn (hartnn):

cot^2 x is given , just plug in values to get csc^2 x!

hartnn (hartnn):

whats 1+ 5/7 =.... ?

OpenStudy (anonymous):

i got 1.71

hartnn (hartnn):

don't use calculator, try to get answer in fraction only...

hartnn (hartnn):

or, you can seee which option comes out to be 1.71......

OpenStudy (anonymous):

i still dont understand

hartnn (hartnn):

1+5/7 = (7+5)/7 = ... ?

OpenStudy (anonymous):

it would be 6/7/=12/7

hartnn (hartnn):

it would be 12/7, yes.

OpenStudy (anonymous):

how did you get 6/7

hartnn (hartnn):

i never got 6/7 i got, \(1+\dfrac{5}{7}= \dfrac{7+5}{7}=\dfrac{12}{7}\)

OpenStudy (anonymous):

and how did you get the 1 there.sorry if im askin so much questions im trying to understand.when i read the question sometimes i dont know how to start them or dnt know what they want to do when it comes to math/

hartnn (hartnn):

1 comes from this relation, \(\huge \csc^2x = 1+ \cot^2 x\)

hartnn (hartnn):

thats the standard relation, which you should've have memorized ...so that you can use here...